Grade 11General Physics

the postion x of a particle moving along x-axis varies with time t as x = Asin(αt) where a and α are positive constants . the acceleration a of the partcle varies with position (x) as
1)a=Ax
2)a= – α2 x
3)a = Aαx
4)a = α2 xA

anirban deb
12 Years agoGrade 11

2 Answers

bharat bajaj
12 Years ago
x = Asin(at)
v = dx/dt = A.a.cos(at)
Acceleration = dv/dt = -A.a^2. sin(at)
Acceleration = -A.a^2.sin(at)
Acceleration = -a^2.x
Hence, 2) is the answerTHanks
Bharat BajajIIT Delhi
askiitians faculty
Arun Kumar
12 Years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty